dominicky:

A triangular prism has a triangular base with dimensions of 3” by 4” by 5” and a volume of 210in^3. What is the surface area of the triangularity prism?

QuestionCoveBot:

No, your incorrect.

QuestionCoveBot:

Your finding the surface area. What you said is the volume of the prism.

rp:

oh

QuestionCoveBot:

So, the 3, 4, 5 can be any side. Base side a, base side b, base side c, and height.

rp:

i used (Perimeter of the base × Length of the prism) + (2 × Base Area) = (S1 + S2+ S3)L + bh

QuestionCoveBot:

So two bases are the same length.

QuestionCoveBot:

Wait, nevermind.

QuestionCoveBot:

I didn't read the whole question. The whole base is 3,4,5 inches.

QuestionCoveBot:

We need to find height.

QuestionCoveBot:

Our values are; Volume - 210^3 Bases - 3,4,5

QuestionCoveBot:

So, we add the basses, 3,4,5 to get 12.

QuestionCoveBot:

Then we have to do some calculations with 12 and the volume (210^3).

rp:

wait i know how to get it the height is three times the volume, divided by the base's area.

QuestionCoveBot:

\[h=\frac{ A _{lat} }{ (a+b+c) }\]

rp:

The formula for the volume of a pyramid is V = (bh) / 3, where b is the area of the base, and h is the pyramid's height. Re-work the equation to solve for h: h = 3V / b.

QuestionCoveBot:

No, RP, your actually wrong. With how it has to be done.

QuestionCoveBot:

I did the calculations that way and it gave me a wrong answer.

rp:

oh thats just how i remember it

QuestionCoveBot:

\[h=\frac{ A _{lat} }{ (3+4+5) }\] \[h=A _{lat}(12)\] \[h=210^3(12)\] \[h=420\div12 = ?\]

QuestionCoveBot:

So what is 420/12 ?

QuestionCoveBot:

My bad, 432, not 420.

QuestionCoveBot:

So 432 divided by 12 is... what, @dominicky

dominicky:

36

rp:

yea good job

QuestionCoveBot:

Correct. So that is your height. I switched things up on accident and did the surface area divided by base and not height times base, my bad. But your final answer for surface area is 432in^2.

rp:

@dominicky you should give him best response

dominicky:

I did thank you so much

QuestionCoveBot:

You're welcome!